Estimation and Confidence Intervals

The Greater Pittsburgh Area Chamber of Commerce wants to estimate the mean time workers who are employed in the downtown area spend getting to work. A sample of 15 workers reveals the following number of minutes spent traveling.

29, 38, 38, 33, 38, 21, 45, 34, 40, 37, 37, 42, 30, 29, 35

Requirements:

i) Develop a 90%, 95% and 98 percent confidence interval for the population mean?

ii) Compare and interpret the result?


1
Expert's answer
2021-06-16T09:28:17-0400
"29, 38, 38, 33, 38, 21, 45, 34, 40, 37, 37, 42, 30, 29, 35"

"n=15"

Sample mean


"\bar{x}= \frac{\displaystyle\sum_{i=1}^nx_i}{n} =\frac{1}{15}
(29+38+38+33+38+21"




"+45+34+40+37+37+42+30+29+35)"

"=\dfrac{526}{15}\approx35.066667"

"Var(x)=s^2= \frac{\displaystyle\sum_{i=1}^n(x_i-\bar{x})^2}{n-1}"

"=\frac{1}{15-1}((29-\dfrac{526}{15})^2+(38-\dfrac{526}{15})^2"

"+(38-\dfrac{526}{15})^2+(33-\dfrac{526}{15})^2+(38-\dfrac{526}{15})^2"

"+(21-\dfrac{526}{15})^2+(45-\dfrac{526}{15})^2+(34-\dfrac{526}{15})^2"


"+(40-\dfrac{526}{15})^2+(37-\dfrac{526}{15})^2+(37-\dfrac{526}{15})^2"

"+(42-\dfrac{526}{15})^2+(38-\dfrac{526}{15})^2+(29-\dfrac{526}{15})^2"




"+(35-\dfrac{526}{15})^2\approx36.2095238"

"s=\sqrt{s^2}\approx6.017435"

i)

90%

The critical value for "\alpha=0.1" and "df=n-1=15-1=14" degrees of freedom is "t_c=1.76131." The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times \dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times \dfrac{s}{\sqrt{n}})"

"=(35.066667-1.76131\times \dfrac{6.017435}{\sqrt{15}},"


"35.066667+1.76131\times \dfrac{6.017435}{\sqrt{15}}













)"

"\approx(32.33013, 37.80321)"

Therefore, based on the data provided, the 90% confidence interval for the population mean is "32.33013<\mu<37.80321," which indicates that we are 90% confident that the true population mean "\mu" is contained by the interval "(32.33013, 37.80321)."


95%

The critical value for "\alpha=0.05" and "df=n-1=15-1=14" degrees of freedom is "t_c= 2.144787." The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times \dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times \dfrac{s}{\sqrt{n}})"

"=(35.066667- 2.144787\times \dfrac{6.017435}{\sqrt{15}},"




"35.066667+ 2.144787\times \dfrac{6.017435}{\sqrt{15}}













)"

"\approx(31.73432, 38.39901)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "31.73432<\mu<38.39901," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(31.73432, 38.39901)."



98%

The critical value for "\alpha=0.02" and "df=n-1=15-1=14" degrees of freedom is "t_c= 2.624494." The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times \dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times \dfrac{s}{\sqrt{n}})"

"=(35.066667- 2.624494\times \dfrac{6.017435}{\sqrt{15}},"




"35.066667+ 2.624494\times \dfrac{6.017435}{\sqrt{15}}













)"

"\approx(30.98900, 39.14433)"

Therefore, based on the data provided, the 98% confidence interval for the population mean is "30.98900<\mu<39.14433," which indicates that we are 98% confident that the true population mean "\mu" is contained by the interval "(30.98900, 39.14433)."


ii)

The 90% confidence interval is "(32.33013, 37.80321)."

The 95% confidence interval is "(31.73432, 38.39901)."

The 98% confidence interval is "(30.98900, 39.14433)."


The 95% confidence interval is wider than the 90% confidence interval.

The 98% confidence interval is wider than the 95% confidence interval and wider than the 90% confidence interval.


Increasing the confidence level increases the error bound, making the confidence interval wider.

Decreasing the confidence level decreases the error bound, making the confidence interval narrower.